OpenStudy (anonymous):

What value of y makes the equation true? 5/-8=y/24 A -15 B -10 C 10 D 15

OpenStudy (texaschic101):

what you have here is a proportion. -5/8 = y/24 cross multiply 8y = (-5)(24) 8y = -120 y = -120/8 y = - 15 check.. -5/8 = -15/24 cross multiply (8)(-15) = (-5)(24) -120 = -120 (correct) so y is -15

OpenStudy (anonymous):

Thank You!

OpenStudy (texaschic101):

glad to help :)